The table below shows the national income statistics for a country.

Items / RM (Million)

Import / RM 6700

Export / RM 8500

Subsidies / RM 480

Consumption / RM 3400

Indirect taxes/ RM 860

Depreciation/ RM 520

Government Expenditure/ RM 4700

Investment / RM 6900

Factor income received abroad/ RM 2010

Corporate tax / RM 620

Factor income paid to abroad/ RM 1800

From the data, calculate:

Gross Domestic Product at market price

Gross National Product at market price

Gross National Product at factor cost

Gross Domestic Product at factor cost

National Income


1
Expert's answer
2022-01-31T10:08:48-0500

(I) Gross Domestic Product at market price


GDPMP = Total national income + sales taxes + Depreciation + factor income received from abroad


= (3400+4700+6900+(8500 - 6700)) + 620 + 520 + 2010

= RM 19950


(II) Gross National Product at market price

GNPMP = GDPMP - factor income paid to abroad


= 19950 - 1800

= RM 18150


(III) Gross National Product at factor cost


GNPFC = GNPMP - net indirect tax


= 18150 - (860 - 480)

= RM 17770


(IV) Gross Domestic Product at factor cost


GDPFC = Gross Value Added at factor cost


GVA = GDPMP + subsidies - taxes on a product


= 19950 + 480 - 860

= RM 19570


(V) National Income


NI = C(consumption)+I(investment)+G(government expenditure)+NX(export - import)

= 3400 + 4700 + 6900 + (8500 - 6700)

= RM 16800